jakfishman:

5(2x − 3) = 5 Part A: How many solutions does this equation have? Part B: What are the solutions to this equation? Show your work.

CocoCammron:

10x -15 = 5 10x = 20 x = 2 only one soultion so it would be x = 2

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CocoCammron 10x -15 = 5 10x = 20 x = 2 only one soultion so it would be x = 2 \(\color{#0cbb34}{\text{End of Quote}}\) brainly

supie:

The first thing you have to do is distribute , \(5(2x−3)=5\) \((5)(2x)+(5)(−3)=5\) \(10x+−15=5\) \(10x−15=5\) Then we have to do \(+15\) to both of the sides (`-15` and we need the opposite) \(10x−15+15=5+15\) \(10x=20\) Then we divide by 10, to both of the sides (it says `10x`) \(\LARGE\frac{10x}{10}=\frac{20}{10}\) \(\frac{10x}{10}\) cancels so all you have left is \(\frac{20}{10}\) Simplify it, and then you'll have your answer.

CocoCammron:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle \(\color{#0cbb34}{\text{Originally Posted by}}\) @CocoCammron 10x -15 = 5 10x = 20 x = 2 only one soultion so it would be x = 2 \(\color{#0cbb34}{\text{End of Quote}}\) brainly \(\color{#0cbb34}{\text{End of Quote}}\) OK first of all I put this in my own wordssssssss

xxDeppressionxx:

Distribute 5(2𝑥−3)=5 10𝑥−15=5 Add 15 to both sides of the equation 10𝑥−15=5 10𝑥−15+15=5+15 add the numbers 10𝑥=20 Divide both sides of the equation by the same term 10𝑥=20 10𝑥/10=20/10 Cancel terms that are in both the numerator and denominator Divide the numbers 𝑥=2 Solution 𝑥=2

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CocoCammron \(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle \(\color{#0cbb34}{\text{Originally Posted by}}\) @CocoCammron 10x -15 = 5 10x = 20 x = 2 only one soultion so it would be x = 2 \(\color{#0cbb34}{\text{End of Quote}}\) brainly \(\color{#0cbb34}{\text{End of Quote}}\) OK first of all I put this in my own wordssssssss \(\color{#0cbb34}{\text{End of Quote}}\) The whole first part was copied word for word but ok..

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle \(\color{#0cbb34}{\text{Originally Posted by}}\) @CocoCammron \(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle \(\color{#0cbb34}{\text{Originally Posted by}}\) @CocoCammron 10x -15 = 5 10x = 20 x = 2 only one soultion so it would be x = 2 \(\color{#0cbb34}{\text{End of Quote}}\) brainly \(\color{#0cbb34}{\text{End of Quote}}\) OK first of all I put this in my own wordssssssss \(\color{#0cbb34}{\text{End of Quote}}\) The whole first part was copied word for word but ok.. \(\color{#0cbb34}{\text{End of Quote}}\) You're still not allowed to do that js

XioGonz:

O.o

jaedengome:

do this in your dms plz

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jaedengome do this in your dms plz \(\color{#0cbb34}{\text{End of Quote}}\) we aren't getting off topic

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jaedengome do this in your dms plz \(\color{#0cbb34}{\text{End of Quote}}\) We're talking about the work Why would be talk about solving the problem in the DM's? Now lets shut up before this gets off-topic and turns into spam.

jakfishman:

thx guys for your help.

supie:

np